OpenStudy (anonymous):

Can someone help me with this Algebra 2?!!!? Its a simple question about an equation..

OpenStudy (anonymous):

Solve the system y = x + 2 and y = -x - 3 by using graph paper or graphing technology. (2 , 4) (-1 , -2) (-2.5 , -0.5) (3.5 , -4.5) I am just wondering if since in the second equation there is a -x, if i have to change any positive x to a negative and a negative would just stay the same??

OpenStudy (anonymous):

Im thinking it would be (-2.5, -0.5) ?

OpenStudy (anonymous):

You would be correct.

OpenStudy (anonymous):

Here you set the equations equal to each other and solve for x. Then you take that result and plug it into either equation to obtain the corresponding y-value.

OpenStudy (anonymous):

Thank you, Ive been staring at this problem for awhile!